OpenStudy (anonymous):

Help ASAP please!! Indicate the equation of the given line in standard form. The line that contains the diagonal of a rectangle with vertices P(2, -2), Q(2, -4), R(7, -2), and S(7, -4) and has a negative slope

OpenStudy (anonymous):

The diagonal could be PS or QR. But we are asked with the -ve slope. Thus \[Slope_{(PS)}=\frac{ -2-(-4) }{ 2-7 }=\frac{ -2 }{ 5 }\]and\[Slope_{(QR)}=\frac{ -4-(-2) }{ 2-7 }=\frac{ 2 }{ 5 }\]since you are asked a line with the -ve slope, we take PS. Thus\[\frac{ y-(-2) }{ x-2 }=\frac{ -2 }{ 5 }\]Then wright it in standard form.

OpenStudy (anonymous):

Thanks! I've got it now!